• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dhamilton

    Log entry time 00:12:01 on November18,2008

    Entry number 249449

    keyword=Swing shift summary

    Swing Shift Summary

    David Hamilton (SL)
    Mark Dalton (TO)
    Diana Parno (3rd)

    16:00 - Shift started with Eugene setting up for the Moeller measurement.
    16:35 - Moeller measurement started.
    19:20 - Moeller measurement finished (see Eugene's entry). Shirish started tuning the beam through the Compton chicane started.
    21:40 - Finished tuning through the Compton. Got back to nominal BPM positions and requested a harp scan.
    22:20 - Target motion problems. Called Yi, JP and Heidi.
    22:45 - Tried to reboot target IOC. No effect. Hall moved to conrolled access to allow JP and Alex to power cycle the motor controllers for the target.
    00:00 - Shift end with hall still at controlled access. Heidi will come to investigate further the problem with target motion.